The de broglie wavelength, associated ith a neutron at teprature 27 degree celcius, is equals to? Also please tell me that i have seen some formulas of thermal wavelenth (in termas of 3mkT, 2mkT, mkT). which one is correct. i am confuse to use them.
1
Expert's answer
2020-06-03T12:02:05-0400

The expression for the de Broglie wavelength associated with a neutron a temperature T is

"\lambda = \frac{h}{\sqrt{3mkT}}"

Here "h=6.626*10^{-34} J.s" is planck's constant.

"m=1.67*10^{-27} kg" is mass of neutron.

"k=1.38*10^{-23} J\/K" is Boltzmann constant.

"T=27^o=300 K" is temperature.

On substituting all above values,

"\lambda = \frac{6.626*10^{-34} J.s}{\sqrt{3(1.67*10^{-27} kg)(1.38*10^{-23} J/K)(300 K)}}"

"=1.45*10^{-10} m =0.145 nm"